mr. mohan travels to his work place first by bus and then by metro. the average speeds of bus and metro are30/hr and 60/hr respectively. the distances covered by bus and metro are in the ratio of1:6. if mr. mohan reaches office in one hour, how far is” his office

Grade:10

1 Answers

Harsh Patodia IIT Roorkee
8 years ago
Distance by bus be x
Distance by metro 6x
Time by bus= x/30
Time by metro= 6x/60
According to question x/30 + 6x/60 = 1
x/30 + x/10 =1
Solve to x= 7.5
Distance is 7x= 7*7.5= 52.5km